OpenStudy (anonymous):

Algebra 2 Word Problem: Set up using 3 variables & 3 equations: "Heather downloaded some television shows. A sitcom uses 0.3gb of memory, a drama .06gb, and a talk show 0.6gb. She downloaded 7 programs totaling 3.6gb. There were twice as many episodes of the drama as the sitcom." A: Write a system of equations for the number of episodes of each type of show. B: How many episodes of each show did she download?

OpenStudy (anonymous):

let: s= #sitcoms d= #drama t= #talk shows so 1 equation can be: s + d + t = 7 try to get the other two....

OpenStudy (anonymous):

that's the easy one. I got that one

OpenStudy (anonymous):

eq 2: 0.3s+0.6d+0.6t=3.6

OpenStudy (anonymous):

I'm not sure about eq3

OpenStudy (anonymous):

eq3 is from this statement: "There were twice as many episodes of the drama as the sitcom."

OpenStudy (anonymous):

2d -s=t?

OpenStudy (anonymous):

no... there should be no "t" for this equation...

OpenStudy (anonymous):

I'm not sure what to do then

OpenStudy (anonymous):

"There were twice as many episodes of the drama as the sitcom." means if I multiply 2*(#sitcoms) that will give me the number of dramas.... does that sound right to you?

OpenStudy (anonymous):

d = 2s

OpenStudy (anonymous):

ie, 2*s = d ??

OpenStudy (anonymous):

yes... :)

OpenStudy (anonymous):

I just got it before you explained it. Ok let me figure out

OpenStudy (anonymous):

notice you can rewrite that as 2s - d = 0

OpenStudy (anonymous):

OK I GOT IT. s=2, d=4, t=1

OpenStudy (anonymous):

Thank you so very much for you help with this. I'm sure I will be posting more questions. I am sick and missed school so I'm catching up work to get ready for midterm

OpenStudy (anonymous):

wait... is that the correct solution?

OpenStudy (anonymous):

oopss.. my bad... that is the correct solution.... :)

OpenStudy (anonymous):

good luck on your midterm.... yw... :)